OpenStudy (anonymous):

if x=1 , 2 , 3, 4, 5 y=7 , 16, 28 , 43, 62 whats the equation, please

OpenStudy (ajprincess):

Let us assume the equation is y=mx+c. Plugging in the first values 7=m*1+c 7=m+c 7-m=c Plugging in the second values 16=m*2+c 16=2m+c 16-2m=c 7-m=16-2m solve for m. Then substitute the value of m in 7-m=c to find c. After finding both m and c check if other values given satisfy ur equation.

OpenStudy (ash2326):

It's not linear, actually a quadratic

OpenStudy (ajprincess):

ya u r right. @ash2326. I see tht other values given do not satisfy the eqn

hartnn (hartnn):

@ash2326 , how did u get quadratic ? if last value was 61, then it would have been quadratic....i think this is exponential or something.....

hartnn (hartnn):

nopes, its 2nd differences are 3,3,4....so higher differences are not constant, thats why it cannot be x^(something)...it might be (something)^x @ajprincess u were telling ?

hartnn (hartnn):

i mean y=a^x+c

OpenStudy (ajprincess):

I tried with the quadratic form bt did nt succeed. maybe u r right. I am not gud with finding exponential eqns. thanx a lot for clearing that up @hartnn

hartnn (hartnn):

u would have got a quadratic equation if last value was 61.... 2nd differences would have been constant,3,3,3 and it would have been polynomial of degree 2